OpenStudy (anonymous):

Solve: 6a^2 + 11a - 10 = 0 a. {-2/5, 3/2} b. {2/5, 2/3} c. {5/2, 2/3} d. {-5/2, 2/3}

OpenStudy (lalaly):

(2a+5)(3a-2) a=-5/2 a=2/3 so d

OpenStudy (anonymous):

Thanks ma'am. :DD

OpenStudy (lalaly):

:):) ur welcome